# Toward architecting self-coding information systems

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Automatic programmingccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software design engineeringccs: Software and its engineering Software design tradeoffs
@@ 2. Self-coding information systems
We define a self-coding information system as an information system whose software part can implement functional requirements or address quality requirements by generating the source code of at least one of its subsystems and using the generated source code at runtime. At the core of the definition is the capability of these systems to generate (or regenerate) their source code. This idea builds on top of existing definitions that hint at autonomy, such as autonomous systems (Adler, 2023), context-aware systems (Schilit et al., 1994), autonomic computing (Kephart and Chess, 2003) (in this case, the authors elaborate on a broad vision of self-* systems that includes autonomos modification of software, but in terms of applying software patches, not self-coding) and self-adaptive systems (see e.g., (Cheng et al., 2008), where the authors do mention adaptation that changes the source code but through dynamic source-code weaving, not actual code generation). While our definition does not prescribe how these systems achieve their self-coding capabilities, agents based on language models (especially those tuned for coding tasks) are nowadays the most straightforward option; therefore, we assume it in this paper.

### v2

@@ Abstract.
copyright: nonejournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ICSE’26 Companion. Please note that this document is a preliminary version; 2026; Accepted at ICSE’26 Track “Software architecture BoF”isbn: XXX-X-XXXX-XXXX-X/XXXX/XXccs: Software and its engineering Automatic programmingccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software design engineeringccs: Software and its engineering Software design tradeoffs
@@ 2. Self-coding information systems
We define a self-coding information system as an information system whose software part can implement functional requirements or address quality requirements by generating the source code of at least one of its subsystems and using the generated source code at runtime. At the core of the definition is the capability of these systems to generate (or regenerate) their source code. This idea builds on top of existing definitions that hint at autonomy, such as autonomous systems (Adler, 2023), context-aware systems (Schilit et al., 1994), autonomic computing (Kephart and Chess, 2003) (in this case, the authors elaborate on a broad vision of self-* systems that includes autonomous modification of software, but in terms of applying software patches, not self-coding) and self-adaptive systems (see e.g., (Cheng et al., 2008), where the authors do mention adaptation that changes the source code but through dynamic source-code weaving, not actual code generation). While our definition does not prescribe how these systems achieve their self-coding capabilities, agents based on language models (especially those tuned for coding tasks) are nowadays the most straightforward option; therefore, we assume it in this paper.